DO NOT USE EXCEL PLEASE .I)A certain car costs €37,500 in Greece and C$42,300 in Canada. If the current exchange rate is C$1.08/€, what is the profit an arbitrageur could earn per car (in terms of Canadian dollars) according to absolutePPP? Round intermediate steps to four decimals and your final answer to two decimals. Use only numeric characters when entering your response.

II)IFE theory relies on the validity of relative PURCHASING POWER PARITY.TRUE OR FALSE

III)Absolute PPP suggests that a given basket of goods should have the same price anywhere in the world.

Iv)Which of the following is a valid criticism of relative PPP?

A

It is only valid if the law of one price is valid.

B

Relative PPP cannot be tested.

C

It ignores factors that may be relevant in forecasting exchange rate fluctuations.

D

None of the above.
Answer #1

according to absolute PPP= exchange rate(C$/Euro)= purchasing price in C$/Purchasing price in Euro= 43000/37500= 1.128

however, the current exchange rate = 1.08

Therefore, there is an arbitrage profit of 1.128-1.08= 0.048C$/Euro

for 37500 Euros, the arbitrage profit is 0.048X 37500= 1800 Euros. This arbitrage profit is calculated assuming there are not frictional costs like transaction/ taxes etc.

2. True

3. A- the relative PPP is derived from Law of one price and therefore have same assumptions.